OpenStudy (anonymous):

please help me i give medals Write the equation 2x + 7y - 4 = 0 in normal form?

OpenStudy (anonymous):

you have to put it in y = mx+b format so you would get y by itself

OpenStudy (anonymous):

no it's like sort form and stuff

OpenStudy (anonymous):

square root*

OpenStudy (ajspeller):

Normal form ??? Never heard that. Do you mean standard form? slope intercept form ?

OpenStudy (anonymous):

m not sure .. my answer choices are: 2x/sqrt(9) + 7y/sqrt(9) - 4/sqrt9) = 0 2x/sqrt(13) + 7y/sqrt(13) - 4/sqrt(13) = 0 2x/sqrt(53) + 7y/sqrt(53) - 4/sqrt(53) = 0 2x/sqrt(69) + 7y/sqrt(69) - 4/sqrt(69) = 0